What is the electric flux when a uniform electric field of 9000 N/C passes through a flat square area of 2m2 at a 300 angle?


1
Expert's answer
2021-06-28T13:39:04-0400

"\Phi=E\cdot A\cdot\cos\alpha=9000\cdot2\cdot\cos(90°-30°)=9000\ (N\cdot m^2/C)" . Answer